OpenStudy (anonymous):

find the slope and the y intercept of the line. whos equation is given y=\[\frac{ 2 }{ 5 }x +6\]

OpenStudy (anonymous):

slope and y-intercept form of the st line is given by y=mx +c comapring y=2/5 x +6 with y=mx+c we have m=2/5 (slope) and c=6 (y-intercept)